CWRU Warner and Swasey Observatory

Visitors
Home
News
Gallery
Reference
Weather

Science Resources
NSRT User Interface
Software
Data Archive
Download Data




Nassau Station

CWRU Nassau Station Robotic Telescope

Public open houses at the telescope site

Telescope status reports.

User News


  • 24 Feb 2003

  • 9/04/02

  • 7/02/02

  • 6/17/02

    • To some extent we are back up and running in a limited mode. There is a personal observing project that is eating most of the available time right now but we'll do public proposals as time allows during the summer shortened nights.

    • The baffle and new secondary mount are completely installed.

    • We expect new software fixing some telescope control problems to be delivered next week.

    • We have discovered a problem with the mirror coating which may force us to re-finish it again.

  • 5/9/02

    • Summer is nearly here in Cleveland and we hope to have the telescope back online soon. The new secondary assembly is in (see the pictures here) but we discovered that we need to put all or part of the sky baffle which we removed back into the telescope.

    • This is the start of open house season visit us if you are in the area.

    • The NSRT Archive has been brought up to date, including all data older than one year old.

  • 4/9/02

    • Why has it been quiet around here? Because our telescope is in pieces, that's why. An explaination of our latest project is here.

    • We have announced the dates for this summer's public open houses.

  • 1/28/02

    • Added some spectacular color pictures of M82 and the Great Orion Nebula in picture gallery 18 and gallery 19.

  • 1/26/02

    • We are experiencing trouble with our FLI camera which is currently being used as our main imager. When the ambient temperature in the telescope dome approaches zero celsius (freezing) the shutter stops working properly.

    • The telescope is scheduled to go down for maintence within the next week for an unknown number of weeks. We are replacing the mount for the secondary mirror with one of sturdier construction that is less susceptible to flexure.

  • 12/12/01

    • Added new images to gallery 18 in the image gallery

    • Wow! Two clear nights in a row. That is rare for winter in Cleveland, OH.

  • 12/11/01

    • Added a new image to gallery 17 in the image gallery

    • Did the first real cold weather test of the telescope last night. Several systems behaved badly. Still need to work out some kinks.

  • 12/3/01

    • Added a new image to gallery 17 in the image gallery

  • 11/17/01

    • Added two new images to gallery 17 in the image gallery

  • 10/16/01

    • An update in the news for the sake of just an update.

    • Our new camera from microluminetics is still not here (it is over one year late now)

    • We had and resolved a problen with moisture in our FLI camera. We were getting ice crystals on the chip when we cooled it but we sent it back to the manufacturer for a retro-fit to fix the problem.

    • The upgraded encoders on the telescope have greatly improved our ability to accurately point the telescope.

    • The auto-dome control still needs work (waiting for word from the contractor).

    • Much progress has been made on the homegrown CCD camera control software as well as the scheduling and data pipeline off the backend of the telescope.

    • We are currently setup to operate in spectroscopic mode. We think we're at the point where we can try really using the spectrograph.

    • Winter is coming on fast here in northeast Ohio and that means fewer clear nights.

    • On the one-year rule, a majority of the data has moved from the Download Area to the Archive.

    • Cosmetic improvements have been made to the Archive access tools.

  • 04/05/01

    • We wanted to be clear on why we are currently not observing with the NSRT. Here are the reasons: 

      • There is a scattered light problem in the telescope that keeps us from being able to flat-field any images. A good illustration of this problem is the picture we took of the Owl Nebula that shows hazy glow around the edges. Previously we had been unaffected by this problem because the camera field of view was smaller, but that changed with the new FLICam. Correcting the problem requires construction of new baffling inside the telescope tube which is on the drawing board and could be installed next as soon as we get the materials needed.

      • We are currently waiting for our new direct imaging camera to arrive from MicoLuminetics.

      • There is more software development that is needed to do any test work with the spectrograph.

  • 03/20/01

    • New rules are being written concerning maximum exposure times on bright objects. We will start enforcing these rules immediately. Proposals in the queue that violate the rules will be altered or thrown out regardless of when that proposal was submitted.

  • 03/19/01

  • 03/17/01

  • 03/16/01

    • Public open houses updated

    • The telescope has been collimated and the new camera software is up and running.

    • The telescope control software is not working and is keeping us from observing.

  • 02/16/01

  • 02/13/01

    • The good news is our mirrors have been re-alluminized and returned to us. The bad news is we still need to clean the interior of the telescope tube, re-mount the mirrors, and collimate the telescope. So still no observing until further notice.

  • 01/16/01

    • January 10th the primary and secondary mirrors were removed from the telescope and will shortly be sent to Pittsburgh, PA for re-alumnizing.

    • The telescope did not make an observation in the month of December or the first half of January due to clouds and extreme cold. The telescope does not operate when the outdoor temperature falls below 15 degrees Fahrenheit because of the hazard that at this temperature the oil in the bearings will congeal and the telescope mechanics will seize up.

  • 12/13/00

  • 11/06/00

  • 11/02/00

    • Added a new feature to the proposal submission forms to alert the user if there is data for the requested object they are requesting in the archive.

    • Added new pictures of the planetary nebulae NGC 2202 and NGC 2371 to picture gallery 15 and the NSRT gallery of planetary nebula.

  • 10/30/00

  • 10/06/00

    • The download area has been updated. Proposals observed since May 2000 now have a file named "index.txt" in each directory which contains a useful information on the FITS files in that directory.

  • 09/30/00

  • 09/28/00

  • 09/27/00

    • Added two new color pictures to the galleries. One of the Blue Snowball nebula (NCG 7662) is in picture gallery 14. And the other of the Saturn Nebula is included with some other pictures taken previouly in picture gallery 5.

  • 09/21/00

    • Added a new color picture of planetary nebula NGC 7027 to picture gallery 14.

    • Announced CWRU open house at the telescope for October 20th

    • Fixed a problem in the download area CGI scripts which caused the script to forget your entered password. Please remember to report problems you have to robotmail@grendel.cwru.edu.

    • Been doing lots of observing. There is a big log jam of proposals in the very early morning we are trying to break up.

    • Had a problem with moisture in our pixcel CCD camera used for direct imaging. It should be fixed. Data taken on Septmber 5th and 6th may have been slightly effected.

    • The new CCD camera for the direct imager and guider are on their way. We expect to start tests with the autoguider system as soon as next week.

  • 08/24/00

    • The tailpiece has been mounted back on the telescope and we are back in bussiness.

  • 08/10/00

    • The telescope is going down. Normal operations will resume after August 20th.

    • People requesting pictures of Pluto should be aware that Pluto has moved too far west in the evening sky to be accessible to the telescope. If your request for Pluto is in the queue and has not yet been fulfulled, it will be another 9 months until we get a chance to do it. Check to see if one of the many pictures we have taken of Pluto that are in the archive meet your needs.

  • 07/20/00

  • 07/10/00

  • 07/06/00

    • Added a picture to Gallery 13.

    • Telescope equipment upgrades are in progess. Expect a new status report soon. The new encoders are looking good. There may be a delay to the spectrograph construction schedule.

  • 06/30/00

  • 06/21/00

  • 06/20/00

  • 06/19/00

  • 06/10/00

  • 05/31/00

  • 05/26/00

  • 05/16/00

    • Updated the download area.

    • Added a script to mail a lost password to a user

    • Good news is that the backend pipeline for the data coming off the telescope is pretty much finished and debuged. A new camera that covers a wider field of view in the direct imaging plane is on order. Bad news is that the telesocpe pointing and control software still has some problems. Hopefully these will get resolved with more clear weather and debugging. We're also having some problems with bugs literally in the optical system. We've had a problem with wasps and yellow jackets getting down into the camera assembly and dying on top of the filters or camera window. We're working on a solution for that as well.

  • 05/04/00

  • 04/02/00

    • Added new simpler proposal submission form in the proposal submission area. Consider the new "snap shot" form the 1040EZ of proposal submission forms.

    • Added a graphical link to the K-12 Teacher Workshops on the main page and user interface.

  • 03/31/00

    • Telescope dome repairs are being done right now. We had a load berring wheel that moves the dome door open and closed that had ceased to turn and was ground flat. When that is replaced we should be ready to run at full tilt again.

    • Added announcements for Summer 2000 public nights at the telescope station.

    • Expanded the tutorial on spectra and spectroscopy in the reference section.

  • 02/04/00

    • The PERL scripts are back online.

  • 02/03/00

    • WE'RE BACK! Well, kind of. None of the PERL scripts that access the databases are working. Be reassured that we lost no data, you just can't get at it over the web right now.

  • 01/27/00

    • Got the web server back online. Only limited parts will be available for atleast a week.

  • 01/22/00

    • Had a disk crash on the server. Did not loose any NSRT data but server is down indefinitely.

  • 01/03/00

  • 12/15/99

    • Updated the color picture of M15 in gallery 5.

    • Updated the picture of Stephan's Quintet in gallery 8.

  • 12/14/99

  • 12/09/99

  • 12/02/99

  • 11/24/99

  • 11/19/99

    • Added the data archive to access older telescope data.

    • New pretty pictures in image gallery #10 since last update.

    • Upgrades are currently in progess on the telescope hardware and we hope to have them completed this weekend.

  • 11/10/99

  • 11/08/99

    • Updated the download area.

    • One week until the telscope goes down for upgrades on November 15th.

  • 11/05/99

  • 11/01/99

  • 10/30/99

  • 10/13/99

    • Updated the download area. Available for immediat public consumption are some new crab nebula images in the "Crab" directory in the public download area.

    • New pretty pictures in the gallery. Look for the Crab Nebula mosaic and color composite of the Little Dumbell Nebula (M76) in gallery #8.

  • 10/05/99

    • Not much activity lately mainly for the reason that we have students from our ASTR 306 (Astronomical Techniques) class using most of the available telescope time to do class projects.

    • Some hardware upgrades are coming in November that should put us a lot closer to running in robotic mode 24/7

  • 09/17/99 

    • Added an announcement for a CWRU community open house at the telescope on the evening of October 8th.

    • Added a mosaic image of Stephan's Quintet to the image galleries in 8.

    • New stuff in the public part of the download area, as of Sept 15.

    • Added graphs of dialy temperature fluctuations and the polaris monitor output to the weather and polaris monitor pages respectively. They've been there on and off for the past week but they are now there to stay.

  • 09/14/99 

  • 09/08/99 

    • Made some changes to the server which for all intents and purposes should be mostly transparent. Please report any bad links you find to robotmail@grendel.cwru.edu

  • 09/06/99 

    • Added many images to the image galleries.

    • New stuff in the public part of the download area.

    • Changes have been made to the proposal submission form to allow request of dark time only exposures (when the moon is not up) and post processing of FITS data before user pickup.

  • 08/30/99 

    • Added many images to the image galleries, including color composite images of the Ring Nebula, Mars, and Uranus.

    • New stuff in the public part of the download area.

Go to the Nassau Station Home Page

Go to the CWRU Nassau Station Robotic Telescope online interface



comments
Last modified July 2, 2002
Case Western Reserve University